OpenStudy (anonymous):

Let f(x)=ax^3+bx^2+cx+d and suppose that r is a root of the equation f(x)=0. Show that r-h is a root of f(x+h)=0.

OpenStudy (rational):

we have below by factor/remainder theorem : `if f(k) = 0, then "k" is a root` say, g(x) = f(x+h) = a(x+h)^3+b(x+h)^2+c(x+h)+d g(r-h) = ?

OpenStudy (anonymous):

g(r-h)=f(r)... Thus, it equals 0??

OpenStudy (anonymous):

@rational ?

OpenStudy (rational):

Exactly ! thats the proof

OpenStudy (anonymous):

@rational Thanks and.... what is the difference between the factor and remainder theorem if you mind answering?

OpenStudy (rational):

thats a very good question !

OpenStudy (rational):

remainder theorem speaks first

OpenStudy (rational):

it talks about the remainder in long division of a "polynomian" by a "linear polynomial" : \[\large \dfrac{P(x)}{x-a}\]

OpenStudy (rational):

the remainder of above long division is same as P(a)

OpenStudy (rational):

thats remainder theorem ^^

OpenStudy (rational):

is that clear ?

OpenStudy (anonymous):

Yes, I think so (: Thank you

OpenStudy (rational):

we're not done with factor theorem yet

OpenStudy (anonymous):

@rational It's okay, I read the definition in the textbook and I remembered it! (: Summer makes me forget a lot ;o Also, will you be willing to maybe solve the last two parts of this problem?

OpenStudy (rational):

good :) factor theorem says this : if \(\large (x-r)\) is a factor of \(\large P(x)\), then \(\large P(r) = 0\) and if \(\large P(r) = 0\), then \(\large (x-r)\) is a factor of \(\large P(x)\)

OpenStudy (rational):

yeah wil try ask

OpenStudy (anonymous):

Thank you (: Show that -r is a root of the equation f(-x)=0

OpenStudy (rational):

use the same trick

OpenStudy (rational):

say, g(x) = f(-x) = a(-x)^3+b(-x)^2+c(-x)+d g(-r) = ?

OpenStudy (rational):

say, g(x) = f(-x) = a(-x)^3+b(-x)^2+c(-x)+d g(-r) = f(--r) = a(--r)^3+b(--r)^2+c(--r)+d = ar^3+br^2+cr+d = 0

OpenStudy (anonymous):

Show that kr is a root of the equation f(x/k)=0

OpenStudy (anonymous):

Oh and thank you (:

OpenStudy (anonymous):

@rational I got it. Thanks